Calculate the net electric field at point A for the following charge distribution. 


q1 northern most point charge of -5x10-5


q2 Eastern most point charge of -3.0x10-5


point a the western most point


point a to q1 is 5.0m


point a to q2 is 5.0m


1
Expert's answer
2021-03-24T20:04:51-0400

Answer

Electric field due to q"_1" On A

"E_1=\frac{kq_1}{r_1^2}\\=\frac{9\times10^9\times5\times10^{-5}}{5^2}\\=1.8\times10^4 V/m"

With direction

"E_1=1.27\times10^4 (-i-j)V/m"

Electric field due to q2 A

"E_2=\frac{kq_2}{r_2^2}\\=\frac{9\times10^9\times3\times10^{-5}}{5^2}\\=0.72\times10^4V/m"

With direction

"E_2=0.72\times10^4(-j)V/m"

So net electric field On Point A

"E=E_1+E_2"

="1.27\times10^4 (-i) +1.99\times10^4( -j)V/m\\=-[1.27i+1.99j]\times10^4 V/m"